# Thermodynamic Entropic Uncertainty Relation

### v1

@@ Abstract
where $\Delta S_{m}$ is the entropy increase in the surrounding medium. Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics. Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{m}$, where $\Delta S_{m}$ represents the increase in entropy in the environment. Then, Eq. (14) can be expressed by $\Delta S_{m}\geq\ln 2-\Lambda[P(\Phi)]$. When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq. (14) is formally identical to the Landauer principle. However, note that the right side of Eq. (17) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution. This contrast with Eq. (14), where the system is assumed to be steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the system state.
Example.—To illustrate Eq. (14), consider a one-dimensional biased random walk as shown in Fig. 2(a). We consider a discrete time model, where the transition probabilities for moving right and left are denoted as $w_{+}$ and $w_{-}$, respectively, with the constraint $w_{+}+w_{-}=1$. The biased random walk plays an important role in biological modeling. For example, a biased random walk can model molecular motors, such as F1-ATPase [31] (Fig. 2(b)). F1-ATPase is an enzyme that can rotate in discrete steps of $32\pi/$ 3 radians in the clockwise or counterclockwise direction [32]. When rotating counterclockwise, F1-ATPase hydrolyzes ATP, while when rotating clockwise, F1-ATPase synthesizes ATP. The rotation in both directions is stochastic, which means that it occurs randomly. The counter-clockwise rotation can be considered as moving in the positive direction in the biased random walk, where the motor has a higher probability of rotating in the counter-clockwise direction.
In the biased random walk, we consider the observable $\Phi(\Gamma)$, which takes either “$+$” or “--”. After the $L$ steps, the observable $\Phi$ is determined by comparing the final position to the initial position. If the final position is greater than the initial position, $\Phi$ is assigned the value “$+$”. Conversely, if the final position is less than or equal to the initial position, $\Phi$ is assigned the value “--”. When the number of steps is odd, there is no possibility of $\Phi(\Gamma)=0$. When considering the aforementioned molecular motor, this observable corresponds to checking whether the motor rotates in the counterclockwise direction or not. For details, see Appendix C.
Figure 3: Results of the biased random walk. The sum of the entropy production and the symmetry entropy, $\Sigma+\Lambda[P(\Phi)]$, is plotted as a function of $w_{+}$ with the solid line. The dotted line denotes $\ln 2$, which is the lower bound of $\Sigma+\Lambda[P(\Phi)]$ in Eq. (14). Each of these terms, $\Sigma$ and $\Lambda[P(\Phi)]$, is plotted by the dashed line and the dot-dashed line, respectively. Here, we employ $L=5$.
@@
We calculate quantities, such as $\Sigma$ and $\Lambda[P(\Phi)]$, as a function of $w_{+}$ in Fig. 3. Note that, for the binary case where $\Phi(\Gamma)$ takes two values, $\Lambda[P(\Phi)]=H[P(\Phi)]$. We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$. In Fig. 3, $\Sigma+\Lambda[P(\Phi)]$, the sum of the entropy production $\Sigma$ and the symmetry entropy $\Lambda[P(\Phi)]$, is represented by the solid line, where the dotted line indicates $\ln 2$. As can be seen in the figure, $\Sigma+\Lambda[P(\Phi)]$ is always above $\ln 2$, confirming the validity of Eq. (14). To examine how each of these components, $\Sigma$ and $\Lambda[P(\Phi)]$, behave as a function of $w_{+}$, $\Sigma$ and $\Lambda[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively. Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $w_{+}$ ranges from $0.5$ to $0.65$. On the other hand, the symmetry entropy of the observable $\Lambda[P(\Phi)]$ is decreasing as a function of $w_{+}$, and except for $w_{+}=0.5$, it is below $\ln 2$. From these results, we can conclude that the inequality $\Sigma+\Lambda[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+\Lambda[P(\Phi)]$.
@@
Conclusion.—In this study, we established a thermodynamic entropic uncertainty relation that links entropy production and the Shannon entropy of the observables. Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics. The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution. We validated this relation using a biased random walk model. This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems. One direction of expansion is towards quantum systems. In recent years, the thermodynamic uncertainty relations in quantum systems [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 14, 43] have garnered significant attention. In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes. This direction presents future challenges.
@@
A.1 Discrete case
In this section, we derive Eq. (16). Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
$P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi\neq 0\\ P(\Phi=0)&\phi=0\end{cases}.$ (18)
@@ A.1 Discrete case
where we abbreviated $P(\phi)=P(\Phi=\phi)$. Then we compute $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
@@
Equation (22) proves Eq. (11) in the main text.
A.2 Continuous case
Next, we consider the continuous case. Basically, the derivation is the same as in the discrete case except that the summation should be replaced by the integration.
Assume $P(\Phi)$ represents a probability density that is differentiable for all $\Phi$. For the continuous case, the probability density $P(|\Phi|)$ is defined by
@@ A.2 Continuous case
We do not consider $P(|\Phi|=0)$, because such measure is 00 for the smooth probability density. When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case. The continuous entropy is given by
$H[P(|\Phi|)]$
[…]
Again, we calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
$H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$
[…]
Using Eq. (25), we prove
@@
Appendix B Proof of Eq. (13)
@@
where we used $H[P(|\Phi|)|P(\Phi)]=0$. Equation (29) proves the first inequality part of Eq. (27). Note that, for the continuous case, this proof does not work, as the continuous conditional entropy $H[P(\Phi)|P(|\Phi|)]$ may take negative values. However, we can show that the first inequality part also holds for the continuous case by a direct calculation.
@@
where we again used the Jensen inequality. Equation (30) proves the second inequality part of Eq. (27). For the continuous case, by following the same argument by replacing the summation with the integration, the following relation holds:
$H[P(|\Phi|)]-H[P(\Phi)]$ $\geq-\ln 2.$ (31)
Appendix C Details of example
We show details of the biased random walk example. The probability distribution at position $x$ after $L$ steps is given by
$p(x)=\begin{cases}{\displaystyle\binom{L}{\frac{L+x}{2}}p^{\frac{L+x}{2}}(1-p)^{\frac{L-x}{2}}}&L+x\in\mathrm{even}\\ 0&\mathrm{otherwise}\end{cases}.$ (32)
Using $p(x)$ in Eq. (32), for odd $L$, the probability of $\Phi$ is calculated by $P(\Phi=\mathrm{decrease})=\sum_{x=-L}^{-1}p(x)$ and $P(\Phi=\mathrm{increase})=\sum_{x=1}^{L}p(x)$, . which can be used to calculate the Shannon entropy of $P(\Phi)$. Moreover, under the steady-state condition, the entropy production $\Sigma$ up to $L$ steps is given by
$\Sigma=L(w_{+}-w_{-})\ln\frac{w_{+}}{w_{-}}.$ (33)

### v3

@@ Abstract
where $\Delta S_{\mathrm{m}}$ is the entropy increase in the surrounding medium. Note that the Landauer principle can be considered as a specific case of the second law of thermodynamics [28]. Given the assumption that the system is in steady state, the change in system entropy $\Delta S$ is zero; consequently, $\Sigma=\Delta S_{\mathrm{m}}$. Then, Eq. (10) can be expressed by $\Delta S_{\mathrm{m}}\geq\ln 2-\Lambda[P(\Phi)]$. When the probability distribution of $P(\Phi)$ ends up with a totally asymmetric distribution, i.e., $\Lambda[P(\Phi)]=0$, Eq. (10) is formally identical to the Landauer principle. However, note that the right side of Eq. (16) arises from a reduction in the Shannon entropy within the system’s state, implying that the Landauer principle is relevant when the system state changes over the time evolution. This contrasts with Eq. (10), where the system is assumed to be in steady state, and $\ln 2$ on its right-hand side arises from the asymmetry of the observable, not the state of the system.
Example.—We demonstrate an application of the thermodynamic entropic uncertainty relation. Decision-making, a fundamental aspect of information processing, is crucial in diverse fields ranging from biology to information science. To illustrate this, we consider the diffusion decision model [29, 30, 31], which describes decision-making as a stochastic process of evidence accumulation. Let $\Gamma$ be a stochastic trajectory of the decision model within the time interval $[0,\tau]$ and $\Phi(\Gamma)$ be a function of $\Gamma$. In the diffusion decision model, a decision occurs when the accumulated evidence during $[0,\tau]$ surpasses thresholds. We assume that if the evidence is greater than or equal to $\theta$, then $\Phi=+1$ (‘yes”‘ class); if it is less than or equal to $-\theta$, then $\Phi=-1$ (“no” class); and in all other cases, $\Phi=0$ (neutral class). Figure 3(a) describes an example of the decision process. Here, $+1,-1,0$ are simply labels, so we can assign other labels. Specifically, let us consider the following Langevin dynamics for modeling the evidence accumulation often employed in the literature [29, 30, 31]:
$\frac{dx}{dt}=f+\sqrt{2A}\xi(t),$ (17)
where $f\in\mathbb{R}$ is a constant force, $\xi(t)$ is the white Gaussian noise with the correlation $\braket{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$, and $A$ is the noise intensity. We define the accumulated evidence as $y(\tau)=\int_{0}^{\tau}\partial_{t}x(t)dt$ and the observable is $\Phi(\Gamma)=1$, $-1$, and 00 when $y(\tau)\geq\theta$, $y(\tau)\leq-\theta$, and $-\theta<y(\tau)<\theta$, respectively. Here, our interest is in the relationship between the entropy of the selected decision $H[P(\Phi)]$ and the entropy production of the Langevin equation. Since $\Phi(\Gamma)$ satisfies the time-reversal antisymmetric condition [Eq. (7)], it satisfies Eqs. (14) and (15), implying that greater entropy production is necessary for lower decision entropy $H[P(\Phi)]$ under the steady-state condition. In decision making, it is well known that there exists a speed-accuracy tradeoff [32, 33], responding faster often reduces accuracy, while achieving higher accuracy lengthens decision time. Through the thermodynamic entropic uncertainty relation, we show quantitatively that another tradeoff, an energy-accuracy tradeoff, holds in the decision model.
@@
We perform numerical calculations for the diffusion decision model [Eq. (17)] to confirm Eq. (15) [26]. We calculate quantities, such as $\Sigma$ and $H[P(\Phi)]$, as a function of $f$ in Fig. 3 for different threshold settings (a) $\theta=0$ and (b) $\theta=0.5$. We consider the steady-state condition and calculate the probability distribution $P(\Phi)$ and $\Sigma$. In Fig. 3, $\Sigma+H[P(\Phi)]$ is represented by the solid line, where the dotted line indicates (a) $\ln 2$ or (b) $(1-P(0))\ln 2$. As shown in Figs. 3(a) and (b), $\Sigma+H[P(\Phi)]$ is always above $\ln 2$ or $(1-P(0))\ln 2$, respectively, confirming the validity of Eq. (15). It is clear that in the case of $\theta=0$, the values of $\Sigma+H[P(\Phi)]$ and the lower bound are closer than in the case of $\theta=0.5$. To examine how each of these components, $\Sigma$ and $H[P(\Phi)]$, behave as a function of $f$, $\Sigma$ and $H[P(\Phi)]$ are represented by the dashed line and the dot-dashed line, respectively. Regarding entropy production $\Sigma$, it falls below $\ln 2$ in the region where $f$ ranges from $0.0$ to $0.25$ in Fig. 3(a). On the other hand, the entropy of the observable $H[P(\Phi)]$ is decreasing as a function of $f$, and except for $f=0.0$, it is below $\ln 2$. From these results, we can conclude that the inequality $\Sigma+H[P(\Phi)]\geq\ln 2$ holds only when considering the sum, $\Sigma+H[P(\Phi)]$, that is, $\Sigma$ or $H[P(\Phi)]$ alone falls below $\ln 2$.
@@
Conclusion.—In this study, we established the thermodynamic entropic uncertainty relation that links the entropy production and the Shannon entropy of the observable. Our findings extend conventional thermodynamic uncertainty relations by incorporating measures based on entropy, highlighting the role of Shannon entropy of observables in stochastic thermodynamics. The derived inequality formalizes a fundamental trade-off between entropy production and the asymmetry of the observable distribution. Through the application to the diffusion decision model, we revealed that accurate decision necessarily require higher entropy production. This framework provides a deeper understanding of nonequilibrium thermodynamics and expands the application of entropy-based uncertainty relations in stochastic systems. We have considered the Kullback-Leibler divergence with respect to trajectories $\Gamma$, but it is also possible to start from the divergence between different quantities (see the supplementary material [26]). Another direction of expansion is towards quantum systems. In recent years, the thermodynamic uncertainty relations in quantum systems [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] have garnered significant attention. In particular, the thermodynamic uncertainty relations within the framework of continuous measurement are closely related to those in classical stochastic processes. This direction presents future challenges.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP23K24915.
@@ Acknowledgements.
Let us introduce the absolute random variable of $\Phi$, which is expressed by $|\Phi|$. The probability distribution of $|\Phi|$ is given by
$P(|\Phi|=\phi)=\begin{cases}P(\Phi=\phi)+P(\Phi=-\phi)&\phi>0\\ P(\Phi=0)&\phi=0\end{cases}$ (18)
@@
where we abbreviated $P(\phi)=P(\Phi=\phi)$.
Using the monotonicity of Kullback-Leibler divergence [Eq. (5)] and the time-reversal property of $\Phi$ [Eq. (7)], we have
$\Sigma$ $\geq D[P(\Phi)\|P(-\Phi)]$
$=-H[P(\Phi)]+C[P(\Phi),P(-\Phi)].$ (20)
Here, the cross entropy is evaluated as $C[P(\Phi),P(-\Phi)]=-\sum_{\phi}P(\Phi=\phi)\ln P(\Phi=-\phi)$. Since the cross entropy term is no less than the entropy, it is non-negative, $C[P(\Phi),P(-\Phi)]\geq H[P(\Phi)]\geq 0$. An important observation is that $C[P(\Phi),P(-\Phi)]$ is even bounded from below by a positive term. We calculate $H[P(|\Phi|)]-C[P(\Phi),P(-\Phi)]$ as follows:
@@
Substituting Eq. (23) into Eq. (20), Equation (23) proves Eqs. (10) and (14) in the main text.
Next, we consider the continuous case. Basically, the derivation is the same as in the discrete case except that summation is replaced by integration. Assume $P(\Phi)$ represents a probability density that is differentiable for all $\Phi$. The Shannon entropy is defined by
[…]
For the continuous case, the probability density $P(|\Phi|)$ is defined by
@@
We do not consider $P(|\Phi|=0)$, because such measure is 00 for the smooth probability density. When $P(\Phi)$ includes the contribution of the delta function in $\Phi=0$, this is not the case. Following the same procedure as the discrete case, we obtain
@@
Appendix B Symmetry entropy and Jensen-Shannon divergence
The Jensen-Shannon divergence is defined by
$\mathrm{JS}[P(X)\|P(Y)]$ $\equiv\frac{1}{2}D\left[P(X)\|\frac{P(X)+P(Y)}{2}\right]$
$+\frac{1}{2}D\left[P(Y)\|\frac{P(X)+P(Y)}{2}\right],$ (27)
which satisfies
$0\leq\mathrm{JS}[P(X)\|P(Y)]\leq\ln 2$ (28)
Using the Jensen-Shannon divergence, we have
$\mathrm{JS}[P(\Phi)\|P(-\Phi)]$
$=-\sum_{\phi}\frac{P(\phi)+P(-\phi)}{2}\ln\frac{P(\phi)+P(-\phi)}{2}$
$+\frac{1}{2}\sum_{\phi}P(\phi)\ln P(\phi)+\frac{1}{2}\sum_{\phi}P(-\phi)\ln P(-\phi)$
$=\ln 2-\sum_{\phi}P(\phi)\ln\left[P(\phi)+P(-\phi)\right]+\sum_{\phi}P(\phi)\ln P(\phi)$
$=\ln 2+H[P(|\Phi|)]-P(0)\ln 2-H[P(\Phi)]$
$=\ln 2-P(0)\ln 2-\Lambda[P(\Phi)],$ (29)
where we used Eq. (19). Equation (29) is Eq. (11) in the main text.
Appendix C Proof of Eq. (9)
@@ Appendix B Symmetry entropy and Jensen-Shannon divergence
where we used $H[P(|\Phi|)|P(\Phi)]=0$. Equation (33) proves the first inequality part of Eq. (30). Note that, for the continuous case, this proof does not work, as the continuous conditional entropy $H[P(\Phi)|P(|\Phi|)]$ may take negative values. However, we can show that the first inequality part also holds for the continuous case by a direct calculation. Specifically, $H[P(\Phi)]-H[P(|\Phi|)]$ is
$H[P(\Phi)]-H[P(|\Phi|)]$
[…]
$=\int_{0}^{\infty}d\phi\,\Bigl{[}[P(\phi)+P(-\phi)]\ln\left[P(\phi)+P(-\phi)\right]$
$-P(\phi)\ln P(\phi)-P(-\phi)\ln P(-\phi)\Bigr{]}.$ (34)
Because $(a+b)\ln(a+b)-a\ln a-b\ln b>0$ for $a>0$ and $b>0$, $H[P(\Phi)]-H[P(|\Phi|)]\geq 0$ is proved.
@@
where we again used the Jensen inequality. Equation (35) proves the second inequality part of Eq. (30).
By adopting the formulation based on the Jensen-Shannon divergence [Eq. (11)], the second part of the inequality in Eq. (30) can be directly derived from Eq. (28). However, note that the first inequality part of Eq. (30) cannot be obtained from Eq. (28).